OpenStudy (anonymous):

A ladder of length 2x+1 feet is positioned against a wall such that the bottom is x−1 feet away from a wall. The distance between the floor and the top of the ladder is 2x feet. Find the length, in feet, of the ladder.

OpenStudy (anonymous):

(2x+1) is the length of ladder which is the hypotenuse 2x is one leg and x-1 is the other leg (2x)^2+(x-1)^2= (2x+1)^2 4x^2+x^2-2x+1=4x^2+4x+1 x^2-6x=0 x^2=6x x=6 Or x=0 x cannot be 0 5,12,13

OpenStudy (anonymous):

What is the 5,12, and 13 at the bottom?

OpenStudy (anonymous):

@mikaylabean

OpenStudy (anonymous):

I see that you got that work from algebra.com, i just don't see where they get 5,12, and 13.

OpenStudy (anonymous):

Okay. I see their error. One second.

OpenStudy (anonymous):

Nevermind, I got it. That was right. Thank you!

OpenStudy (anonymous):

Mmhm!